What is the y-component of a 100 N force that is exerted at 60N to the horizontal
1
Expert's answer
2019-03-05T09:40:24-0500

The y-component of a force


"F_y=\sqrt{F^2-F_x^2}=\sqrt{100^2-60^2}=80 N"

Answer: 80 N